Circular No. 8508

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005bg IN MCG +03-31-93 R. Quimby, P. Mondol, F. Castro, M. Sellers, P. Hoeflich, and J. C.

Wheeler, University of Texas; and C. Gerardy, Imperial College, report the discovery of a supernova in unfiltered CCD images taken with the 0.45-m ROTSE-IIIb telescope at McDonald Observatory on Mar. 28.17 (with the new object at mag ∼ 16.2) and Apr. 2.16 UT (mag ∼ 16.1). SN 2005bg is located at α = 12^{h}17^{m}17^{s}.18, δ = +16^{o}22′17′′.6 (equinox 2000.0), which is 0′′.6 north and 0′′.4 east of the core of MCG +03-31-93. The new object was found by subtracting a co-addition of images taken between 2004 Dec. 11 and 2005 Jan. 17 (limiting mag ∼ 18.6); a similar subtraction showed nothing at this location on Mar. 20.19 (limiting mag ∼ 17.2). A spectrum (range 430–890 nm) of the new object, obtained on Apr. 7.13 with the 9.2-m Hobby/Eberly Telescope (+ Marcario Low-Resolution Spectrograph) by B. Roman and S. Rostopchin, shows it to be a type-Ia supernova. Assuming the NED host-galaxy recession velocity of 6921 km/s, the expansion velocity derived from the minimum of the Si II 635.5-nm feature is 11000 km/s. The spectrum also shows Mg II 448.1-nm, Fe II 508.3-nm, and S II 545.4and 564.0-nm lines; however, the Ca II infrared triplet is not detected. Additional unfiltered CCD magnitudes for SN 2005bg: Apr. 3.16 UT, 15.8; 5.15, 15.6.

SUPERNOVA 2005bd AND 2005be M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 340–730 nm) of SN 2005bd (cf. IAUC 8505), obtained on Apr. 9.15 UT by H. Landt with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a type-Ia supernova with a spectral-feature age (cf. Riess et al. 1997, A.J. 114, 722) of ∼ 8 (± 2) days after maximum light. The supernova expansion velocity, derived from the minimum of Si II (rest 635.5 nm) and adopting the NED recession velocity of 11119 km/s for the host galaxy, is ∼ 11000 km/s.

M. Salvo, B. Schmidt, and J. Norris, Australian National University (ANU), report that a spectrum (range 360–530 nm) of SN 2005be (IAUC 8506) was taken at the ANU 2.3-m telescope (+ Double-Beam spectrograph) at Siding Spring on Apr. 6.70 UT. Cross-correlation with a library of supernovae spectra (see IAUC 8098) shows SN 2005be to be a type-Ia supernova nearly a week past maximum light, at a redshift z = 0.035. This is the same redshift value found for the host galaxy from the position of the Ca H and K lines in its spectrum.

2005 April 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green